OpenStudy (anonymous):

How do you complete this problem (some people are giving me reciprocals, and it's not right!)? x divided by z

OpenStudy (anonymous):

\[\frac{x}{z}\]

OpenStudy (anonymous):

x*z^-1 then?

OpenStudy (anonymous):

x equala 6 and z equals 1 and a half

OpenStudy (anonymous):

equals**

OpenStudy (anonymous):

4

OpenStudy (anonymous):

How did you write the 1 and a half?

OpenStudy (anonymous):

Oh. So plug in those numbers: \[\frac{6}{\frac{1}{2}}\] Multiply the top and bottom by 2 to get rid of the complex fraction: \[= \frac{6}{\frac{1}{2}} \cdot \frac{2}{2}\]\[= \frac{12}{1} = 12\]

OpenStudy (anonymous):

Oh lame, 1 and a half, not 1 half.

OpenStudy (anonymous):

So you have a 3 on bottom, not a 1, and it's \[\frac{12}{3} = 4\]

OpenStudy (anonymous):

Sorry! I mean't one half! :0(

OpenStudy (anonymous):

Oh, then it's what I said the first time ;)

OpenStudy (anonymous):

Turn the bottom one upside down and multiply it instead

OpenStudy (anonymous):

Sorry, invert the denominator...

OpenStudy (anonymous):

I always hated the method of 'turning' fractions. Multiplying top and bottom by the inner fraction's denominator always made more sense to me.

OpenStudy (anonymous):

Personally I've always found the whole idea of "fractions" a bit off...